OpenStudy (anonymous):

If an Italian restaurant offers 5 different toppings to put on pizzas how many different pizzas can be prepared? NOTE: assume the restaurant also prepares pizza with no toppings.

OpenStudy (raden):

5!

OpenStudy (raden):

do u have the choices ?

OpenStudy (anonymous):

16 31 25 32

jimthompson5910 (jim_thompson5910):

Hint: Number of ways = (number of ways with 5 toppings) + (number of ways with 4 toppings) + (number of ways with 3 toppings) + (number of ways with 2 toppings) + (number of ways with 1 toppings) + (number of ways with 0 toppings)

OpenStudy (anonymous):

16?

jimthompson5910 (jim_thompson5910):

so you're adding 5 C 5 + 5 C 4 + 5 C 3 + 5 C 2 + 5 C 1 + 5 C 0 you might notice that these numbers can be found in pascals triangle and they add up to some power of a whole number

OpenStudy (anonymous):

5+4+3+2+1+1= 16

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

there is 5 C 5 = 1 way to order a pizza with 5 toppings there are 5 C 4 = 5 ways to order a pizza with 4 toppings there are 5 C 3 = 10 ways to order a pizza with 3 toppings etc etc add them all up

OpenStudy (anonymous):

32?

jimthompson5910 (jim_thompson5910):

yep, a short cut is to look at pascals triangle and you'll see a row starting with 1, 5, 10, ... add those numbers up to get 1+5+10+10+5+1 = 32

OpenStudy (anonymous):

awesome thanks!

jimthompson5910 (jim_thompson5910):

an even bigger shortcut is to use the idea that the answer is 2^n where n = number of toppings you can choose from so 2^5 = 32

OpenStudy (anonymous):

oh i see